# Algebraic Connectivity Enhances Hyperedge Specificity in the Alzheimer’s Disease Continuum

### v1

@@ 2.2.2 Hyperedge Modeling
Different methods have been proposed for computing the hyperedge weights ($\mathbf{W}$ matrix). The $\mathbf{W}$ matrix is a fundamental aspect when working with hypergraphs, since via this information different features can be extracted, such as the node degree matrix $\mathbf{D}_{v}$ and the hypergraph similarity matrix $\mathbf{S}$. To this end, in this work, we proposed a novel method of computing the weight of each hyperedge, relying on the algebraic connectivity ($a(\mathcal{G})$, Fiedler value) fiedler1973algebraic . The importance of $a(\mathcal{G})$ resides in the fact that it is a measure of how well connected a graph is de2007old . The $a(\mathcal{G})$ is computed on the subgraph composed of the regions belonging to a specific hyperedge. In detail, the absolute value of the FC matrix defined by the regional time series belonging to the hyperedge was computed, thus generating a subgraph of the brain regions of the hyperedge. The $a(\mathcal{G})$ is defined as the second-smallest eigenvalue extracted from the Laplacian matrix of the subgraph. Due to the limited number of regions involved in each hyperedge (at least three), LU Factorization was used to decompose the Laplacian matrix.
@@ 2.3.1 Baseline Approaches
The $a(\mathcal{G})$ approach was compared via a post-hoc analysis with three common approaches to compute the hyperedge weights, and with the mean correlation value of the Schaefer functional networks (FNs), where each FN was considered as a hyperedge. In detail:
@@
The Gaussian similarity kernel, defined as $K[i,j]=\exp{\frac{-||\mathbf{x}_{i}-\mathbf{x}_{j}||_{2}^{2}}{2\sigma^{2}}}$, was computed across the time series of the regions belonging to the hyperedge, where $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are the $i$-th and $j$-th time series, respectively, and $\sigma$ was set using the heuristic median technique.
@@
The mean correlation was computed across all the pairwise correlations calculated between pairs of time series belonging to the hyperedge.
@@
The L2 norm of the LASSO coefficients in the hyperedge in the form of $\mathbf{w}_{m}=||\boldsymbol{\alpha}_{m}||_{2}^{2}$.
@@
The mean correlation values calculated from the seven FNs of the Schaefer atlas were employed as the hyperedge weights, where the hyperedges were expressed in terms of FNs regions.
@@ 2.3.3 Mediation Analysis
Entorhinal cortex, a small region in medial temporal lobes, is among the earliest sites affected by AD pathology and is considered a key origin point for tau accumulation igarashi2023entorhinal . From this condition, pathological tau spreads to other brain regions via neural connections, contributing to disease progression and cognitive decline. The spread of tau pathology in AD is believed to follow a prion-like mechanism, advancing transneuronally across synaptically and functionally connected regions wang2024characterization . At the same time, tau deposition is known to disrupt FC, suggesting a reciprocal relationship between network integrity and tau progression roemer2025amyloid ; franzmeier2022tau . Motivated by this interplay, we conducted a mediation analysis to explore whether high-order functional features captured by hypergraph-derived hyperedge weights might mediate the link between entorhinal cortex tau burden, as measured by SUVR from tau-PET, and cognitive performance. The same analysis was performed on the hyperedge weights extracted from the Schaefer FNs. The analysis was conducted on a subset of the original cohort, including only participants with available entorhinal cortex tau SUVR values provided by ADNI and corresponding cognitive scores, and considering all diagnostic groups jointly. As a result, five MCI and three AD patients were excluded due to missing data. The following linear regressions modeled the relations between variables:
@@
Figure 3: Hyperedges found statistically relevant in the proposed analysis with their FDR-corrected p-values for the four different approaches.
Figure 4: Boxplots representing the distribution of the weights of the three groups for the statistically significant hyperedges and FNs.
@@ 3.1 Group Differences in Extracted Weights
A non-parametric statistical analysis was performed on the hyperedge weights extracted from all the approaches to assess differences across HC, MCI, and AD subjects. Figure 3 shows the brain regions forming hyperedges that exhibited statistically significant differences across groups. The corresponding FDR-corrected p-values are also reported. Table 2 summarizes the number of regions and their associated FNs for each significant hyperedge. Figure 4 highlights the distribution of these hyperedge weights, stratified by group and weighting method. Results showed that hyperedges 9, 48, 60, and 75, which include regions from the somatomotor, visual, and default mode FNs, exhibited reduced $a(\mathcal{G})$ in AD compared to both HC and MCI. In contrast, hyperedges 22 and 77, composed of regions within the dorsal and ventral attention FNs, showed increased $a(\mathcal{G})$ in AD patients. Similarly, the other approaches highlighted the same patterns with higher values in HC than MCI and AD for the hyperedges 9 and 75.
@@
The weights computed with the $a(\mathcal{G})$ demonstrated greater discriminative power compared to the other three approaches, identifying six hyperedges with significant differences, while the mean Pearson correlation, the L2 norm of the LASSO coefficients, and the Gaussian similarity kernel identified only two significant hyperedges. All four methods highlighted differences across groups for hyperedge 9, which comprises regions within the somatomotor network. Hyperedge 75, composed of four regions from the salience ventral attention network and one region from the somatomotor network, was identified as significant by three of the methods, including the $a(\mathcal{G})$. When examining the FNs hyperedge weights, three out of seven FNs were found to be statistically different: the somatomotor, salience ventral attention, and default mode networks.
Table 2: Information of the statistically significant hyperedges.
Hyperedge Brain regions Functional networks involved 9 3 Somatomotor 22 3 Dorsal Attention (2), Ventral Attention (1) 48 5 Visual (2), Default (3) 59 5 Somatomotor (4), Ventral Attention (1) 60 7 Somatomotor (5), Ventral Attention (2) 75 5 Somatomotor (1), Ventral Attention (4) 77 3 Dorsal Attention (2), Ventral Attention (1)
@@
To further assess pairwise group differences, post-hoc comparisons were performed using the Mann-Whitney U test. Table 3 summarizes the results of this analysis, highlighting the number of hyperedges that exhibited significant differences between each pair of groups (HC vs MCI, HC vs AD, and MCI vs AD).
@@
Table 3: Post-hoc analysis results showing the number of different hyperedges between group pairs.
@@
The post-hoc analysis revealed a consistent ability across all considered approaches to discriminate HC from AD and MCI from AD. On the other hand, no hyperedge weights were able to differentiate the HC and MCI groups, except for the hyperedge 9 identified by the LASSO-based method. Similarly, the somatomotor network of the Schaefer atlas allowed differentiating HC and MCI groups.
@@ 3.2 Mediation Analysis Results
Table 4: Statistically significant results of the mediation analysis with the bootstrap method (significance set at $p<0.05$) examining the relationship across tau entorhinal cortex SUVR, algebraic connectivity values, and five cognitive composite scores. Values represent the coefficients of the regression forming the total, direct, and indirect effects.
@@
Hyperedge Algebraic connectivity - $a(\mathcal{G})$ Memory Executive Language Visual-Spatial ADAS-13 60 Total: -0.479 Direct: -0.466 Indirect: -0.014 Total: -0.214 Direct: -0.200 Indirect: -0.014 Total: 6.246 Direct: 6.076 Indirect: 0.170 75 Total: -0.479 Direct: -0.459 Indirect: -0.021 Total: -0.303 Direct: -0.283 Indirect: -0.019 Total: -0.214 Direct: -0.195 Indirect: -0.019 Total: -0.164 Direct: -0.149 Indirect: -0.015 Total: 6.246 Direct: 5.950 Indirect: 0.296
@@
Table 4 summarizes the results of the mediation analysis performed using $a(\mathcal{G})$. Among the six hyperedges that showed significant group differences, only two revealed significant mediation effects. Specifically, hyperedge 60 showed a partial mediation effect between tau and the memory, language, and ADAS-13 composite scores. Hyperedge 75 revealed a broader significant partial mediation effect, linking the tau accumulation with all the cognitive scores. Across the analysis, the regression coefficients showed a similar pattern in the two hyperedges. In detail, a negative relation was found between tau burden and the five composite cognitive scores, where higher tau values were related to cognitive decline, e.g., lower values in the composite cognitive scores; while a positive effect was shown for the ADAS-13 score, where higher values were related to the disease.
@@
The Gaussian similarity kernel and the mean Pearson correlation found significant mediation analysis for the weights of hyperedge 75 and all the cognitive scores, while the L2 norm of the LASSO coefficients approach found only significant relations in hyperedge 9 and for the memory and ADAS-13 scores. No significant mediation effects were found when using the Schaefer FNs-derived weights as mediator.

### v2

@@ 2.2.2 Hyperedge Modeling
Different methods have been proposed for computing the hyperedge weights ($\mathbf{W}$ matrix). The $\mathbf{W}$ matrix is a fundamental aspect when working with hypergraphs, since via this information different features can be extracted, such as the node degree matrix $\mathbf{D}_{v}$ and the hypergraph similarity matrix $\mathbf{S}$. To this end, in this work, we proposed a novel method of computing the weight of each hyperedge, relying on the algebraic connectivity ($a(\mathcal{G})$, Fiedler value) Fiedler (1973). The importance of $a(\mathcal{G})$ resides in the fact that it is a measure of how well connected a graph is De Abreu (2007). The $a(\mathcal{G})$ is computed on the subgraph composed of the regions belonging to a specific hyperedge. In detail, the absolute value of the FC matrix defined by the regional time series belonging to the hyperedge was computed, thus generating a subgraph of the brain regions of the hyperedge. The $a(\mathcal{G})$ is defined as the second-smallest eigenvalue extracted from the Laplacian matrix of the subgraph. Due to the limited number of regions involved in each hyperedge (at least three), the trace minimization with LU factorization approach was used to decompose the Laplacian matrix, and the tolerance value for computing the eigenvalues was set to 1e101e-10 in order to have numerical stability when calculating the $a(\mathcal{G})$.
@@ 2.3.1 Baseline Approaches
The $a(\mathcal{G})$ approach was compared via a post-hoc analysis with three common approaches to compute the hyperedge weights, and with the mean correlation value of the Schaefer 100 functional networks (FNs), where each FN was considered as a hyperedge. In detail:
@@
[…]
@@
The mean correlation was computed across all pairwise correlations calculated between pairs of time series belonging to the hyperedge.
@@
[…]
@@
The mean correlation values calculated from the seven FNs of the Schaefer 100 atlas were employed as the hyperedge weights, where the hyperedges were expressed in terms of FNs regions.
@@ 2.3.2 Statistical Analysis
For each significant hyperedge, the Cliff’s $\delta$ effect size was computed in absolute value between group pairs and then averaged to obtain a global effect size across the three groups.
@@
2.3.3 Binary Classifications
Three binary classification tasks have been modeled not to outperform the state-of-the-art in AD classification, but to assess the potential of the extracted functional features from the hypergraph to discriminate between the different clinical groups. To this end, the statistically different hyperedge weights derived from each method have been used, separately, as input features for a random forest model to classify, respectively, (i) HC vs AD, (ii) MCI vs AD, and (iii) HC vs MCI. For each task, an initial train-test split (80/20) was performed, using the training set to tune the random forest hyperparameters and train the model using 5-fold cross validation, while retaining the test set as a hold-out set for the test phase. The same train/test sets were used for all methods employed to compute the hyperedge weights, but each random forest was optimized independently. The models’ performance was reported in terms of accuracy, precision, recall, and F1 score for the hold-out test set. Finally, the internal features importance of the random forest models was extracted to understand which hyperedge contributed most to the classification.
@@ 2.3.4 Mediation Analysis
Entorhinal cortex, a small region in the medial temporal lobes, is among the earliest sites affected by AD pathology and is considered a key origin point for tau accumulation Igarashi (2023). From this condition, pathological tau spreads to other brain regions via neural connections, contributing to disease progression and cognitive decline. The spread of tau pathology in AD is believed to follow a prion-like mechanism, advancing transneuronally across synaptically and functionally connected regions Wang and others (2024). At the same time, tau deposition is known to disrupt FC, suggesting a reciprocal relationship between network integrity and tau progression Roemer-Cassiano and others (2025); Franzmeier and others (2022). Motivated by this interplay, we conducted a mediation analysis, with and without covariates (age, sex, years of education, and APOE4), to explore whether high-order functional features captured by hypergraph-derived hyperedge weights of all methods might mediate the link between entorhinal cortex tau burden, as measured by SUVR from tau-PET, and cognitive performance. The same analysis was performed on the weights extracted from the Schaefer 100 FNs, where each FN was considered as a hyperedge. To note, this analysis was conducted on a subset of the original cohort of ADNI-3 and considering all diagnostic groups jointly. More specifically, five MCI and three AD patients did not have the entorhinal cortex tau SUVR values and/or cognitive scores available, and for this reason, they were excluded from the analysis.
The following linear regressions modeled the relations between variables:
@@
Figure 4: Hyperedges found statistically relevant in the proposed analysis with their FDR-corrected p-values for the four different approaches (FDR level of 0.05). The highlighted regions were mainly part of salience/ventral attention, somatomotor, dorsal attention, default-mode, and visual FNs.
Figure 5: Boxplots representing the distribution of the weights of the three groups for the statistically significant hyperedges and FNs. In the figure, the hyperedges are denoted by the letter "e" followed by the corresponding hyperedge number used in this study. For each hyperedge, the Kruskal-Wallis FDR-corrected p-value ($\mathrm{p}<0.05$) and corresponding effect size (ES) are annotated.
@@
2.3.5 External Validation
We tested the robustness of the $a(\mathcal{G})$ considering an atlas with higher resolution. We extracted the fMRI time series of the Schaefer 400 atlas from the fMRI volumes for all the subjects, resulting in 400 regional time series with 192 time points each. The hypergraph backbone was extracted using the same HC-hypergraph individuals employed for defining the hypergraph structure of the Schaefer 100 atlas. Finally, considering the same cohort of HC-analysis, MCI, and AD subjects from the previous analyses, we conducted the statistical analyses described in 2.3.2 for all hyperedge weight computation methods.
We additionally tested the potential of the $a(\mathcal{G})$ on a different cohort of subjects from Phase 2 of ADNI (ADNI-2), who did not overlap with the selected cohort of ADNI-3. A total of 96 individuals were collected, including 21 HC (Age: $75.8\pm 8.1$; Sex: M=9/F=12; MMSE: $28.9\pm 1.7$), 42 MCI (Age: $73.5\pm 8.6$; Sex: M=22/F=20; MMSE: $27.5\pm 2.1$), and 33 AD (Age: $74.6\pm 6.7$; Sex: M=15/F=18; MMSE: $21.3\pm 4.5$). We considered the same hypergraph backbone computed on the HC-hypergraph subjects of ADNI-3 based on the Schaefer 100 atlas, but statistical analyses across the three groups and for post-hoc pairwise tests as described in 2.3.2 were performed on the new cohort.
@@ 3.1 Group Differences in Extracted Weights
The $\mathcal{HG}$ structure was extracted using the HC-hypergraph group for the subsequent analyses. The stability of the resulting $\mathcal{HG}$ structure was evaluated with respect to the length of the time series. Stable results were obtained for fMRI data comprising more than 135 time points. At this lower bound, comparison between $\mathcal{HG}$ structures extracted from signals with 135 and 192 time points yielded a Jaccard index of 0.855 and a correlation of 0.921. Consistently, the number of hyperedges was comparable, with 90 and 96 hyperedges for the 135 and 192 time points signals, respectively.
@@
Based on these results, only the $\mathcal{HG}$ structure computed from the full-length time series was used in the subsequent analysis. A non-parametric statistical analysis was performed on the hyperedge weights extracted from all the approaches to assess differences across HC, MCI, and AD subjects. Figure 4 shows the brain regions forming hyperedges that exhibited statistically significant differences across groups. The corresponding FDR-corrected p-values are also reported. Table 2 summarizes the number of regions and their associated FNs for each significant hyperedge. Figure 5 illustrates the distribution of these hyperedge weights, along with their corresponding FDR-corrected p-values and effect sizes, stratified by group and weighting method. Results showed that hyperedges 9, 48, 60, and 75, which include regions from the somatomotor, visual, and default-mode FNs, exhibited reduced $a(\mathcal{G})$ in AD compared to both HC and MCI. In contrast, hyperedges 22 and 77, composed of regions within the dorsal and salience/ventral attention FNs, showed increased $a(\mathcal{G})$ in AD patients. Similarly, the other approaches highlighted the same patterns with higher values in HC than MCI/AD for hyperedges 9 and 75. Further, the global Cliff’s $\delta$ effect size revealed a small separation effect across the three groups, with the higher values for hyperedge 75.
@@
The weights computed with the $a(\mathcal{G})$ highlighted a larger number of significant hyperedges compared to the other three approaches, identifying six hyperedges with significant differences. The mean Pearson correlation, the L2 norm of the LASSO coefficients, and the Gaussian similarity kernel identified only two significant hyperedges. All four methods highlighted differences across groups for hyperedge 9, which comprises regions within the somatomotor FN. Hyperedge 75, composed of four regions from the salience/ventral attention FN and one region from the somatomotor FN, was identified as significant by three of the methods, including the $a(\mathcal{G})$. When examining the FNs hyperedge weights, three out of seven FNs were found to be statistically different: the somatomotor, salience/ventral attention, and default-mode FNs.
Table 2: List of regions involved in the statistically different hyperedges. For each region, the region name of the Schaefer 100 atlas, the MNI centroid coordinates, the hyperedge containing the region, and the main overlap in percentage with the anatomical regions of the Harvard-Oxford atlas are reported. The FNs are included in the region name (Vis=Visual; SomMot=Somatomotor; DorsAttn=Dorsal attention; SalVentAttn=Salience/Ventral attention; Default=Default-mode), while LH and RH mean left and right hemisphere. Post. and ant. means posterior and anterior division.
Region name MNI Centroid Coordinates Hyperedges Main overlaps (Harvard-Oxford atlas) LH Vis 1 (-26, -34, -18) 48 Parahippocampal Gyrus, post. (62%); Temporal Fusiform Cortex, post. (22%) LH Vis 6 (-12, -66, 6) 48 Intracalcarine Cortex (53%); Lingual Gyrus (11%) LH SomMot 1 (-54, -22, 8) 9 Heschl’s Gyrus (with H1 and H2) (35%); Planum Temporale (32%) LH SomMot 2 (-36, -22, 16) 59 Insular Cortex (31%); Central Opercular Cortex (24%) LH SomMot 3 (-54, -12, 14) 9, 60 Central Opercular Cortex (68%) LH DorsAttn FEF 1 (-26, -4, 60) 22 Superior Frontal Gyrus (33%); Middle Frontal Gyrus (15%) LH SalVentAttn FrOperIns 1 (-42, -2, -6) 75 Insular Cortex (64%) LH SalVentAttn FrOperIns 2 (-38, 12, 6) 75 Frontal Operculum Cortex (59%); Central Operculum Cortex (21%) LH SalVentAttn Med 3 (-6, 4, 62) 22, 77 Supplementary Motor Cortex (49%) LH Default pCunPCC 1 (-12, -56, 12) 48 Precuneous Cortex (44%) LH Default pCunPCC 2 (-6, -52, 34) 48 Precuneous Cortex (50%); Cingulate Gyrus (42%) RH SomMot 1 (52, -16, 6) 9, 59, 60 Heschl’s Gyrus (with H1 and H2) (43%); Planum Temporale (12%) RH SomMot 2 (40, -16, 16) 59, 60 Central Opercular Cortex (41%); Insular Cortex (24%) RH SomMot 3 (56, -4, 12) 59, 60, 75 Central Opercular Cortex (48%) RH SomMot 4 (58, -6, 30) 60 Postcentral Gyrus (39%); Precentral Gyrus (29%) RH DorsAttn FEF 1 (28, -2, 60) 22, 77 Superior Frontal Gyrus (38%); Middle Frontal Gyrus (21%) RH SalVentAttn TempOccPar 2 (60, -26, 28) 60, 75 Supramarginal Gyrus, ant. (49%); Parietal Operculum Cortex (19%) RH SalVentAttn FrOperIns 1 (40, 8, 2) 59, 60, 75 Insular Cortex (50%); Central Opercular Cortex (13%) RH SalVentAttn Med 2 (8, 6, 52) 77 Supplementary Motor Cortex (62%) RH Default pCunPCC 1 (12, -54, 14) 48 Precuneous Cortex (56%)
@@
To further assess pairwise group differences, post-hoc comparisons were performed using the Mann-Whitney test. Table 3 summarizes the results of this analysis, highlighting the number of hyperedges that exhibited significant differences between each pair of groups (HC vs MCI, HC vs AD, and MCI vs AD).
@@
Table 3: Post-hoc analysis showing the number of significant hyperedges between group pairs over the total number of hyperedges found in each method. The approaches could discriminate the HC-AD and MCI-AD groups, but not HC-MCI.
@@
The post-hoc analysis revealed a consistent ability across all considered approaches to discriminate between HC and AD, as well as MCI and AD. On the other hand, no hyperedge weights were able to differentiate the HC and MCI groups, except for hyperedge 9 identified by the LASSO-based method. Similarly, the somatomotor FNs of the Schaefer 100 atlas allowed differentiating HC and MCI groups.
@@
3.2 Classification Results
Table 4: Classification results for the three tasks: (i) HC vs AD (top), (ii) MCI vs AD (center), and (iii) HC vs MCI (bottom).
HC vs AD ACC PRE REC F1 Algebraic connectivity 0.800 0.831 0.800 0.798 Gaussian similarity kernel 0.657 0.656 0.657 0.654 Mean Pearson correlation 0.771 0.771 0.771 0.771 L2 norm LASSO coef. 0.657 0.663 0.657 0.658 MCI vs AD ACC PRE REC F1 Algebraic connectivity 0.750 0.750 0.750 0.691 Gaussian similarity kernel 0.696 0.506 0.696 0.586 Mean Pearson correlation 0.679 0.503 0.679 0.578 L2 norm LASSO coef. 0.714 0.510 0.714 0.595 HC vs MCI ACC PRE REC F1 Algebraic connectivity 0.695 0.687 0.695 0.609 Gaussian similarity kernel 0.695 0.666 0.695 0.644 Mean Pearson correlation 0.644 0.586 0.644 0.593 L2 norm LASSO coef. 0.678 0.645 0.678 0.644
Three classification tasks were addressed, relying on the hyperedge weights as input features for the random forest model for each hyperedge weight computation method. The hyperedge weights extracted using the $a(\mathcal{G})$ showed a more discriminative power in the three tasks, except for the HC vs MCI comparisons, where comparable performance was obtained by the Gaussian similarity kernel approach. In detail, in the first task (HC vs AD), $a(\mathcal{G})$ achieved an accuracy of 0.800, while the Gaussian similarity kernel, mean correlation, and the L2 norm of the LASSO coefficients obtained accuracies of 0.657, 0.771, and 0.657, respectively. For the second task (MCI vs AD), accuracies of 0.750, 0.696, 0.679, and 0.714 were achieved by $a(\mathcal{G})$, Gaussian similarity kernel, mean correlation, and L2 norm of LASSO coefficients, respectively, showing how the $a(\mathcal{G})$ outperformed the other methods. Finally, the model on the third task (HC vs MCI) yielded the same result of 0.695 using the $a(\mathcal{G})$ as well as the Gaussian similarity kernel. In contrast, 0.644 and 0.678 were obtained by mean correlation and L2 norm of the LASSO coefficients, respectively. Table 4 shows the performance of the methods for the three tasks, reporting accuracy, precision, recall, and F1 score.
The importance of the features of the different models across the tasks revealed that hyperedge 75 was the most important for $a(\mathcal{G})$. Similarly, hyperedge 75 was relevant for the Gaussian similarity kernel and mean correlation methods in discriminating MCI from AD and HC from MCI, while hyperedge 9 was relevant for distinguishing the HC group from AD. Finally, hyperedge 59 was the most important in all tasks for the L2 norm of the LASSO coefficients.
@@ 3.3 Mediation Analysis Results
Table 5: Statistically significant results of the mediation analysis with the bootstrap method (significance set at $p<0.05$) examining the relationship across tau entorhinal cortex SUVR, algebraic connectivity values, and five cognitive composite scores, showing both results without (top) and with covariates (age, sex, years of education, and APOE4) (bottom). Values represent the coefficients of the regression forming the total, direct, and indirect effects, with their corresponding confidence intervals [2.5%,97.5%] and p-value (p-val).
@@
Hyperedge Mediation analysis for Algebraic connectivity - $a(\mathcal{G})$ - without covariates Memory Executive Language Visual-Spatial ADAS-13 Coeff p-val Coeff p-val Coeff p-val Coeff p-val Coeff p-val 60 $\mathrm{w_{m}}\sim\mathrm{z}$ -0.108 [-0.211,-0.005] 0.0394 -0.108 [-0.211,-0.005] 0.0394 -0.108 [-0.211,-0.005] 0.0394 $\mathrm{y}\sim\mathrm{w_{m}}$ 0.178 [ 0.096, 0.260] <<0.001 0.148 [ 0.087, 0.210] <<0.001 -2.228 [-3.255,-1.202] <<0.001 Total -0.479 [-0.548,-0.412] <<0.001 -0.214 [-0.274,-0.154] <<0.001 6.246 [ 5.417, 7.076] <<0.001 Direct -0.466 [-0.533,-0.399] <<0.001 -0.200 [-0.259,-0.142] <<0.001 6.076 [ 5.257, 6.896] <<0.001 Indirect -0.014 [-0.034,-0.001] 0.046 -0.014 [-0.031, 0.000] 0.046 0.170 [ 0.008, 0.409] 0.048 75 $\mathrm{w_{m}}\sim\mathrm{z}$ -0.192 [-0.294,-0.090] <<0.001 -0.192 [-0.294,-0.090] <<0.001 -0.192 [-0.294,-0.090] <<0.001 -0.192 [-0.294,-0.090] <<0.001 -0.192 [-0.294,-0.090] <<0.001 $\mathrm{y}\sim\mathrm{w_{m}}$ 0.199 [ 0.117, 0.280] <<0.001 0.158 [ 0.083, 0.233] <<0.001 0.136 [ 0.074, 0.199] <<0.001 0.106 [ 0.041, 0.172] 0.002 -2.684 [-3.699,-1.699] <<0.001 Total -0.479 [-0.548,-0.412] <<0.001 -0.303 [-0.373,-0.233] <<0.001 -0.214 [-0.274,-0.154] <<0.001 -0.164 [-0.228,-0.099] <<0.001 6.246 [ 5.417, 7.076] <<0.001 Direct -0.459 [-0.527,-0.390] <<0.001 -0.283 [-0.354,-0.212] <<0.001 -0.195 [-0.255,-0.135] <<0.001 -0.149 [-0.214,-0.084] <<0.001 5.950 [ 5.119, 6.781] <<0.001 Indirect -0.021 [-0.043,-0.008] <<0.001 -0.019 [-0.042,-0.007] 0.006 -0.019 [-0.036,-0.007] <<0.001 -0.015 [-0.031,-0.005] 0.008 0.296 [ 0.130, 0.559] <<0.001 Hyperedge Mediation analysis for Algebraic connectivity - $a(\mathcal{G})$ - with covariates Memory Executive Language Visual-Spatial ADAS-13 Coeff p-val Coeff p-val Coeff p-val Coeff p-val Coeff p-val 75 $\mathrm{w_{m}}\sim\mathrm{z}$ -0.192 [-0.306,-0.078] <<0.001 -0.192 [-0.306,-0.078] <<0.001 -0.192 [-0.306,-0.078] <<0.001 -0.192 [-0.306,-0.078] <<0.001 -0.192 [-0.306,-0.078] <<0.001 $\mathrm{y}\sim\mathrm{w_{m}}$ 0.137 [ 0.061, 0.212] <<0.001 0.113 [ 0.040, 0.186] 0.002 0.088 [ 0.028, 0.148] 0.004 0.079 [ 0.013, 0.145] 0.019 -1.988 [-2.946,-1.031] <<0.001 Total -0.457 [-0.527,-0.387] <<0.001 -0.269 [-0.346,-0.193] <<0.001 -0.213 [-0.275,-0.148] <<0.001 -0.126 [-0.198,-0.053] <<0.001 5.934 [ 5.048, 6.819] <<0.001 Direct -0.444 [-0.515,-0.373] <<0.001 -0.255 [-0.332,-0.178] <<0.001 -0.201 [-0.265,-0.137] <<0.001 -0.114 [-0.187,-0.041] 0.002 5.725 [ 4.833, 6.617] <<0.001 Indirect -0.013 [-0.033,-0.001] 0.046 -0.014 [-0.037,-0.001] 0.036 -0.011 [-0.027,-0.001] 0.004 -0.012 [-0.031,-0.002] 0.032 0.209 [ 0.059, 0.459] 0.004
@@
Table 5 summarizes the results of the mediation analysis performed using $a(\mathcal{G})$ with and without covariates. Among the six hyperedges that showed significant group differences, only two revealed significant mediation effects without considering the covariates. Specifically, hyperedge 60 showed a partial mediation effect between tau and the memory, language, and ADAS-13 composite scores. Hyperedge 75 revealed a broader significant partial mediation effect, linking the tau accumulation with all the cognitive scores. Across the analysis, the regression coefficients showed a similar pattern in the two hyperedges. In detail, negative relationships were found between tau burden and memory, executive, language, and visual-spatial scores, where higher tau values were associated with cognitive decline, e.g., lower values in the composite cognitive scores. Conversely, a positive effect was observed for the ADAS-13 score, where higher values were associated with the disease. The same analysis was performed, controlling for the covariates. Only the weights of hyperedge 75 showed a significant partial mediation effect with all the scores. The same patterns observed in the regression coefficients for hyperedge 75 without covariates were also found after controlling for age, sex, years of education, and APOE4. In general, we also observed a reduction in the total effects of the mediation analysis when including the covariates, and consequently, the indirect effect was decreased with respect to the mediation analysis without covariates.
@@
The Gaussian similarity kernel and the mean Pearson correlation revealed significant mediation analysis for the weights of hyperedge 75 and all the cognitive scores with/without covariates (except for the visual-spatial score where no mediation effect was found with mean Pearson correlation), while the L2 norm of the LASSO coefficients approach found only significant relations in hyperedge 9 and for the memory (with/without covariates) and ADAS-13 (without covariates) scores. No significant mediation effects were found when using the Schaefer 100 FNs-derived weights as mediators.
3.4 External Validation Results
The statistical analysis described in 2.3.2 for the Schaefer 100 atlas was repeated using the high-resolution Schaefer 400 atlas to assess the robustness of $a(\mathcal{G})$. $a(\mathcal{G})$ identified five significant hyperedges after FDR correction, while the Gaussian similarity kernel, mean correlation, and L2-norm LASSO approaches identified six, zero, and one significant hyperedges, respectively. Although the Gaussian kernel found one additional hyperedge, $a(\mathcal{G})$ showed stronger discriminative power in post-hoc pairwise tests, yielding significant differences (FDR-corrected) for all hyperedges across all group comparisons. Conversely, the Gaussian similarity kernel identified five out of six hyperedges as significantly different for HC vs AD and MCI vs AD, and two out of six hyperedges for HC vs MCI. Furthermore, since the FDR correction was applied to a large number of hyperedges, we also investigated a less stringent significance threshold (p << 0.10). Under this threshold, both $a(\mathcal{G})$ and the Gaussian similarity kernel identified the same number of hyperedges (10). The five additional hyperedges of $a(\mathcal{G})$ showed a trend towards significance. Specifically, their FDR-corrected p-values were close to 0.05, namely 0.054, 0.058, 0.058, 0.061, and 0.081. However, the effect size of the significant hyperedges was computed, and all the methods showed a small effect size across the three groups. Biologically, the five significant hyperedges identified by $a(\mathcal{G})$ were composed of regions mainly corresponding to the salience/ventral attention, default-mode, control, and somatomotor FNs. Instead, the five close-to-significance hyperedges were associated with the visual and dorsal attention FNs. Notably, the dorsal and salience/ventral attention, default-mode, somatomotor, and visual FNs were also relevant when considering the Schaefer 100 atlas.
Similarly, the same analyses described in 2.3.2 were performed using the individuals of ADNI-2. None of the methods detected significant hyperedges after FDR correction. However, using raw p-values, $a(\mathcal{G})$ identified 10 hyperedges, while the Gaussian similarity kernel, the mean correlation, and the L2 norm of the LASSO coefficients detected a lower number (four with both the Gaussian similarity kernel and the mean correlation, and three with the L2 norm of the LASSO coefficients). Considering these hyperedges, in post-hoc pairwise comparisons, $a(\mathcal{G})$ demonstrated the ability to differentiate between group pairs (8/10 significantly different hyperedges for HC vs MCI, 7/10 for HC vs AD, and 2/10 for MCI vs AD). The Gaussian kernel detected mainly HC vs AD differences (0/4 for HC vs MCI, 4/4 for HC vs AD, 2/4 for MCI vs AD), mean correlation showed effects in HC vs MCI and HC vs AD (3/4 for HC vs MCI, 3/4 for HC vs AD, 0/4 for MCI vs AD), and LASSO highlighted changes across all pairs but most strongly in MCI vs AD (2/3 for HC vs MCI, 1/3 for HC vs AD, 3/3 for MCI vs AD). Additionally, two hyperedges found to be statistically different in the ADNI-2 cohort were also identified as significant in ADNI-3, considering $a(\mathcal{G})$. A small effect size was found for all methods across groups. In this case, the hyperedges were composed mainly of regions of the salience/ventral attention, dorsal attention, somatomotor, and default-mode FNs.